OpenStudy (anonymous):

Please help me solve... P and Q are two points at distance x(x>R,radius of the Earth) and 2x respectively from the centre of the Earth.The gravitational potential at P is -800kj/kg.What is the work done to take a 1kg mass from P to Q?

OpenStudy (anonymous):

@Abhisar

OpenStudy (abhisar):

Have you tried it urself ?

OpenStudy (abhisar):

What is the answer given ?

OpenStudy (amriju):

400 i think...

OpenStudy (anonymous):

400kj

OpenStudy (abhisar):

\(\sf \huge U=\frac{-GmM}{r}=\large Potential\times mass ~of~the~object\) Since mass of the object=1, potential in this case will be equal to potential energy. Now, potential at point P is -800kj, potential energy stored in the object of mass 1 kg will be -800kJ. Also potential energy of the same object at point Q will be -800/2=-400kj. Now work done in bringing the object from P to Q will be equal to the energy difference i.e -400-(-800)=400kj

OpenStudy (abhisar):

Got it ?

OpenStudy (anonymous):

Ya...got it...

OpenStudy (abhisar):

Great !